OpenStudy (anonymous):

3x+6+1=8 A. Subtraction Property of Equality B. Combine like terms C. Distributive Property D. Division Property of Equality

OpenStudy (anonymous):

3x+7=8 3x+7-7=8-7 3x=1 3x/3=1/3 x=1/3

OpenStudy (anonymous):

no no i know that but what is it a b c or d? thats the part i dont get

OpenStudy (anonymous):

i think except C every other one is right !

OpenStudy (anonymous):

you can only pick one

OpenStudy (anonymous):

then B

OpenStudy (anonymous):

but its sill pretty confusing

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

well good luck :D

OpenStudy (anonymous):

thanks